Given f(x)=4x+1 and g(x)=x^2+1
Find: (f o g)(x)=
my answer was 4x^2+4 was I correct?

Found 2 solutions by mukhopadhyay, venugopalramana:
Answer by mukhopadhyay(490)

(f o g)(x)=
f (x^2+1) = 4(x^2+1) + 1 = 4x^2 + 5
Your answer was very close (off by 1)

Answer by venugopalramana(3286)

Given f(x)=4x+1 and g(x)=x^2+1
Find: (f o g)(x)=
LET Y = G(X)= X^2+1
F(G(X))=F(Y)=4Y+1 =4(X^2+1)+1=4X^2+4+1=4X^2+5
my answer was 4x^2+4 was I correct?..NO..IT IS 4X^2+5